OpenStudy (anonymous):

PLZ HELP ME!!! Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent $625. Write an equation in the standard form to represent form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. Part B: Write the equation obtained in part A using function notation. Part C: Describe the steps to the graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

I WILL HELP U IF U HELP ME 11. (03.06 LC) Choose the equation below that represents the line that passes through the point (-7, 2) and has a slope of −3. (5 points) y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2) Score: 0 of 5

OpenStudy (anonymous):

Ok give me a sec

OpenStudy (anonymous):

Use the point slope formula Y-y1=m(x-X1)

OpenStudy (anonymous):

So it y-2=-3(x-(-7)) This become positive So the answer is a A- y-2=-3(x+7)

OpenStudy (anonymous):

Can u PLZ help me now